Find the area of a triangle bounded by the x-axis,y-axis and the line 3x - 4y=12
Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

Find the area of a triangle bounded by the x-axis,y-axis and the line 3x - 4y=12
-------------
find the x-intercept and the y-intercept.
Area = (x-int)*(y-int)/2

Answer by Boreal(15235)

Standard form, find the intercepts.
x=0; -4y=12; y= -3
y=0; 3x=12; x=4
The bounds mean the x=axis is 4 units long to the intercept (4,0)
The y-intercept means that there are 3 units between 0 and -3.
This is a 3,4,5 right triangle so the hypotenuse is 5 units, along the line
The base is 4, the height is 3, and the area is (1/2)bh=6 units.
ANSWER:6 UNITS.